### v2

# VURF: A General-purpose Reasoning and Self-refinement Framework for Video Understanding

Abstract
Recent studies have demonstrated the effectiveness of Large Language Models (LLMs) as reasoning modules that can deconstruct complex tasks into more manageable sub-tasks, particularly when applied to visual reasoning tasks for images. In contrast, this paper introduces a Video Understanding and Reasoning Framework (VURF) based on the reasoning power of LLMs. Ours is a novel approach to extend the utility of LLMs in the context of video tasks, leveraging their capacity to generalize from minimal input and output demonstrations within a contextual framework. By presenting LLMs with pairs of instructions and their corresponding high-level programs, we harness their contextual learning capabilities to generate executable visual programs for video understanding. To enhance program’s accuracy and robustness, we implement two important strategies. Firstly, we employ a feedback-generation approach, powered by GPT-3.5, to rectify errors in programs utilizing unsupported functions. Secondly, taking motivation from recent works on self refinement of LLM outputs, we introduce an iterative procedure for improving the quality of the in-context examples by aligning the initial outputs to the outputs that would have been generated had the LLM not been bound by the structure of the in-context examples. Our results on several video-specific tasks, including visual QA, video anticipation, pose estimation and multi-video QA illustrate the efficacy of these enhancements in improving the performance of visual programming approaches for video tasks. https://github.com/ahmad-573/VURF
Keywords:
Video Understanding Reasoning Vision Language
1 Introduction
Figure 1: An Overview of the VURF pipeline: Figure demonstrates how a complex query regarding video editing is broken down in our Video Programming framework to sequentially arrive at the final edited result. Best viewed in zoom.
In recent years, the vision community has developed highly efficient specialized models for various video understanding tasks, including Video Question Answering [4], Action Anticipation [23] and Pose Estimation [58, 31, 70, 21]. Despite such advancements, the video models usually offer isolated visual comprehension capabilities in the form of narrow task-specific models. Such specialized models limited to individual tasks struggle to offer a comprehensive, adaptable, and scalable understanding of videos for complex reasoning. Although the current off-the-shelf models can perform specific well-defined tasks, they require specialized and comprehensive datasets to effectively train dedicated models for each task, which is not feasible for general-purpose complex reasoning problems. Furthermore, individual off-the-shelf vision models for each video-understanding task necessitate a distinct framework with unique model configurations. This problem underscores the requirement for a uniform reasoning framework that offers plug-and-play architecture, capable of leveraging any pre-trained computer vision model, enabling seamless execution of a given task.
To address these challenges, we adopt a strategy of decomposing broad video-understanding tasks into more manageable sub-tasks, each of which can be solved by executing task-specific models, and subsequently consolidating the results. Our framework is motivated by the observation that complex tasks can be effectively solved by executing an intermediate sequence of sub-tasks, collectively working in sequence to solve the more challenging problem. This process of task decomposition necessitates a reasoning module capable of discriminating the necessary steps for task execution. Large Language Models (LLMs) emerge as promising candidates for this role. Recent works on Visual Programming [27] has demonstrated the effectiveness of LLMs in converting complex tasks into smaller, more tractable components that can be addressed by specialized computer vision models [64, 69, 47]. In our study, we demonstrate the utility of such a reasoning module in tackling specific challenges within the domain of video understanding. We also show that building such an approach on top of the existing off-the-shelf video models can significantly enhance task performance.
While LLMs demonstrate competence in serving as reasoning modules, they are not immune to errors and limitations. One notable deficiency is their vulnerability to hallucinations induced by contextual information, without the means to self-correct based on task-agnostic knowledge [54]. To address this concern, we draw inspiration from recent research demonstrating the efficacy of self-refinement processes in enabling LLM’s to enhance their outputs [44, 19], akin to the way humans engage in self-correction. Specifically, we propose a feedback-generation mechanism that evaluates the LLM’s output and employs this feedback to prompt the LLM to refine its output. Our findings substantiate the effectiveness of this approach in elevating the performance of video-based reasoning approaches. Our main contributions are as follows:
  • The first generic visual reasoning framework for video understanding consolidating multiple task-specific, domain-specialized video models to answer any video-related user queries.
  • Using in-context learning, we align LLM behavior for decomposing a given complex task into multiple subtasks easily solvable using existing task-specific video models.
  • Our proposed self-refinement strategy helps avoid errors in the programs (outlining subtask decomposition) generated by the LLM and boosts the performance by iteratively refining the generated program. The proposed framework is shown to boost performance for tasks such as visual question answering for videos in complex reasoning scenarios.
2 Related Work
Video Understanding: Video understanding focuses on teaching machines to understand and analyze visual content. One crucial task is to recognize and localize different actions in the video [46]. Numerous algorithms have been developed to cater the video understanding tasks such as Video Swin Transformer [41], VideoMAE [57], C2D [61], MViT-V2 [38], STGCN++ [71], and ViViT [5] that achieve high accuracy on SOTA datasets [51, 22, 26, 15, 28, 23, 48].
Video understanding tasks facilitate the efficient handling of diverse information modalities [36] and various tasks have been introduced to test the capabilities of the methods for video understanding such as retrieving temporal and spatial information [76] and answering natural language questions from the video i.e., Video Question Answering (VQA) [68, 34]. Other methods such as SeViLA [72] and iVQA [40], adapt Localizer and Answerer for both QA and temporal key-frame localization which are then extended to zero-shot VQA [52, 67, 39]. Another challenging task in video understanding is to localize the starting and ending time of the video segment that corresponds to the input query i.e., video grounding [11, 67, 75]. It can solve various video understanding tasks such as Temporal Action Recognition [12], spatio-temporal video grounding [77], and Action Recognition [9]. Our work on video programming provides a pipeline for various video understanding and reasoning tasks leveraging off-the-shelf SOTA models and the reasoning capabilities of LLMs for each sub-task.
Visual Programming: Visual Programming leverages the capabilities of Language Models (LLMs) as reasoning modules to break down complex vision-understanding tasks into simpler intermediate sub-tasks executed sequentially. Utilizing in-context examples along with prompts can effectively decompose these tasks, enhancing the responses generated by language models. In recent advancements within visual programming, such as VisProg [27], strategies have been proposed to enhance performance in visual tasks. These strategies include increasing the number of in-context examples, substituting high-error modules with recent off-the-shelf models, and fine-tuning instructions to achieve improved performance. Our work, VURF, stands out as the first generic reasoning framework for the video modality. It not only utilizes the self-critique strategy of LLMs but also employs periodic refinement of visual programs. This process further ensures the replacement of invalid functions with a valid module list, addressing misalignment errors effectively.
The concept of Visual Programming can be extended to any family of vision-language tasks such as BERT, and it’s inspired vision-language work [20, 37, 80]. Recently, zero-shot models such as Flamingo have emerged that do not require fine-tuning on new tasks [2]. These body of works solve different vision-understanding tasks such as Visual Question Answering, VLM for few-shot learning [2], Grounding, Segmentation, and Tracking. Moreover, the vision-language model incorporates multiple types of information and is trained using a contrasting approach which has made it possible to adapt to new tasks without the necessity of fine-tuning, allowing for zero-shot adaptation [29]. However, zero-shot models struggle to generalize and outperform SOTA in the unseen data as they lack adaptation on specific tasks [33, 60]. In contrast, VURF, leverages SOTA off-the-shelf models for the downstream tasks by decomposing complex tasks into visual programs that uses SOTA models.
Self Refinement: LLMs demonstrate a special ability to enhance their outputs just like how humans re-evaluate, refine, and reiterate the text that they have initially written. The same LLM when used to identify the potential issues with the output and through the generator, refiner, and feedback provider, it even has the potential to improve responses generated by SOTA LLMs like GPT4 [43]. Other methods include using external tools like search engines to rectify the output [43], detecting hallucinated outputs [25, 18, 79, 24], using natural language feedback [49] to improve the initial generated response.
Leveraging the capabilities of self-debugging, language models have been able to debug their predicted program via few-shot demonstrations [13], use of relevant in-context examples to generate efficient response [59, 56], and have contributed in various domains such as code generation and its applications [73], summarization [7], and program synthesis [32, 30].
While these approaches improve the LLM’s response via recursive-feedback methods, these methods require continuous refinement of any output generated. Our self-refinement approach focuses on the critique and refinement of in-context examples and we show that just a pre-refined set of in-context examples can boost the performance of a visual programming approach (Table 1).
Figure 2: Video Understanding and Reasoning Framework (VURF) pipeline. Top: figure shows the main approach of visual programming for videos with the added self-correction module. Bottom: figure shows the self-refinement module. Our proposed VURF framework allows improved reasoning for video understanding building on the existing individual task-specific models
3 Methodology
In contrast to conventional task-specific models that exhibit limitations in addressing complex reasoning challenges, the Video Understanding and Reasoning Framework (VURF) seeks to utilize the reasoning power of LLMs to deconstruct complex video-related queries into a series of sub-tasks (video programs). By executing these sub-tasks sequentially, we can culminate to arrive at the final response. Moreover, VURF allows seamless integration of new visual models in the plug-and-play manner and also employs the self-critique mechanisms to mitigate LLM’s hallucinations and judgement errors. An overview of our approach is shown in Fig. 2, and we explain the approach in detail below.
3.1 Video Reasoning LLMs
Role of Large Language Models. Large language models (LLMs), exemplified by GPT-3 [6] and GPT-4 [1], have demonstrated an impressive in-context learning capability (ICL) that does not necessitate fine-tuning. ICL aims to extend the understanding of LLMs to novel scenarios using a restricted set of input and output demonstrations within the relevant context. In this work, we leverage GPT-3.5 to generate visual programs that solves video reasoning tasks involving natural language instructions. The LLMs efficiently performs video reasoning tasks by avoiding direct video processing. Instead, it formulates logical sequence flows that decompose complex tasks into simpler sub-tasks. This approach enhances efficiency and allows the model to navigate the intricacies of video-related challenges.
Prompting. We prompt GPT-3.5 with in-context examples which consists of pairs of instruction and the associated programs that the LLM is expected to generate. The programs follow a generic structure where each line of the program includes the name of a module, the module’s input argument names and their values, and an output variable name. As output variables in a specific step are used later for another step they follow a general structure which the LLM learns:
    OUTPUT0=FUNC0(video=VIDEO,...)
    OUTPUT1=FUNC1(arg0=OUTPUT0,...)
Given a set of these pairs and a new instruction, the LLM can generate a new program that follows the same structure and can thus be executed via our program interpreter.
3.2 Self-Refinement
One prominent limitation of a naive visual programming technique like [27] is the proneness of generating inaccurate information influenced by contextual cues, as well as lacking an inherent capacity for autonomous self-correction through task-agnostic knowledge. There are two major issues that arise with the LLM-generated program. Firstly, due to LLM hallucinations, the program might be using some function that is not supported by our interpreter. Secondly, the program may not break down the instruction into sub-tasks in an optimal fashion. We resolve these limitations by the below steps.
Error Correction: To address the issue of a program utilizing a function unsupported by our interpreter, we employ a feedback generation approach, which notably leverages the power of GPT-3.5. We present the program to this module, alongside the available list of functions and their general usage, and inquire if the given program violates these constraints. If discrepancies are identified, the program is subsequently regenerated, incorporating the provided feedback as contextual information. This iterative process ensures error-free execution of the given instruction.
Auto-Refinement of In-Context Examples: The effectiveness of our approach hinges significantly on the quality of in-context examples. Therefore, it is imperative to refine these examples to enhance the module’s performance [42, 43]. To accomplish this, we implement a self-refinement procedure. Initially, given an instruction, $I$, and the initial program generated $P$, we input $I$ to the LLM, prompting it to generate an improved program, $P^{\prime}$, without the inclusion of in-context examples. Subsequently, both $P$ and $P^{\prime}$ are input into the LLM, enabling the generation of a refined program that aligns more closely with the LLM’s reasoning, all while excluding the influence of in-context examples. Replicating this process for $n$ instructions yields a set of $n$ new in-context examples, enhancing the module’s ability to perform tasks effectively. Note that as shown in Fig. 2, the auto self-refinement module can be applied to a single user query to iteratively improve the generated program as well. However, this is inefficient and costly and thus we show that even pre-refining the in-context program-instruction pairs can improve the performance of the VURF (Table 1). A concrete illustration of this process on a single program and instruction pair is provided in Fig. 3.
Figure 3: Auto Self-Refinement demonstrated through an example. Two programs are generated: one is provided with contextual examples, and the other one is not - although certain information is appended to ensure the program’s structural integrity. Subsequently, both programs are input into the Language Model (LLM), with the task of generating a new program that closely aligns with the ideal program but does not use any invalid functions.
Manual Refinement of In-Context Examples: While the auto self-refinement approach enhances contextual relevance, it falls short in augmenting the functionality of the in-context examples. To overcome this limitation, we propose a methodology that involves leveraging the capabilities of the Large Language Model (LLM) to introduce new functions or modules (Fig. 4), thereby elevating the quality of our in-context examples. Our approach entails generating programs from a set of instructions and tasking the LLM with enhancing the program, without imposing constraints on the functions it may introduce. Through careful observation of the resulting program, we can identify and incorporate valuable additions to our list of functions.
Figure 4: Main Modules used by VURF. The red boxes show modules that require a pre-trained model whereas the boxes are modules that require trivial functions.
4 Tasks
Our video understanding and reasoning framework is an effort to provide a comprehensive approach that can be easily adapted to various visual tasks by incorporating an interpreter component for an existing pre-trained SOTA vision model. VURF solves four diverse and comprehensive tasks, including Video Question Answering (VQA), Video Anticipation, Pose Estimation, and Multi-Video VQA. The program works by using LLM as a reasoning module to generate a visual program that consists of subsequent steps that are executed in isolation and the output of each predecessor as fed to the input to the successor.
4.1 Video Question Answering
Figure 5: A qualitative example showing the Program steps in the Multi-Video VQA task. The programs provide a logical decomposition of the original complex tasks.
VQA is an important task in video comprehension that endeavors to connect natural language processing with video comprehension. The objective is to empower models to interpret the content of a video and respond to user queries related to the video content. Effective resolution of such queries often involves breaking down the overarching problem into manageable subproblems.
For instance, consider a video of a man entering a room and then performing some action. If a human is asked, “What does the man do after entering the room?” they would visually inspect the video, locate the man, observe his actions by examining the video and then deduce what the nature of his activities. This example illustrates the critical role of logical reasoning in decomposing complex questions into more manageable sub-queries, suitable for evaluation using off-the-shelf models. The inherent complexity of such tasks makes VQA an ideal candidate for a visual programming approach, given its primary function of decomposing intricate tasks into more manageable components. In the context of the aforementioned example, VURF, as demonstrated in Fig. 7 (Right), initially employs a GROUNDING model to identify an interval where the man enters the room. Subsequently, it would call the TRIMAFTER module to retrieve the relevant part of the video and finally using VQA module, it will engage in visual examination of the video to answer the question “Pick up towel”. This approach makes the task more logical and interpretable, with possible explanations in case an output is wrong.
VURF also extends its functionality to Multi-Video Question Answering tasks which involves synthesizing information from two distinct videos to provide accurate responses to user queries. Our approach simplifies this intricate process by breaking it down into a sequence of steps executed by a language model. For instance, in Fig. 5, when presented with the user query “Are the number of people in the kitchen in both videos the same?”, our approach generates a visual program. This program initiates the TRACK step in both videos to identify and isolate regions containing people. Subsequently, the COUNT step computes the number of people in each video, and a final answer is derived through a comparison of these counts.
Figure 6: Qualitative example showing the program steps of the Pose Estimation task of VURF. We show limited frames due to space constraints. Best viewed in zoom
4.2 Pose Estimation
Pose Estimation identifies the position and orientation of human bodies within images or videos and uses a skeletal model to represent them [78]. This field has drawn significant attention in the research and has been utilized in various domains such as HCI [16], virtual reality [45], clinical use in pediatric disorders [17], and motor assessment in adults [53]. Numerous methods have been employed to solve pose estimation. The general workflow mainly operates in two stages, first, identify and locate the orientation of joints in the human body e.g., by using ‘Human model’ in MMPose [50], neural network architecture in OpenPose [8], and through the 3D-2D Keypoints Matching [10]. The next step is to use a framework or model that inputs this graph skeleton and outputs the estimated pose of the object [3].
Although these model-based approaches in pose estimation frameworks perform well on numerous problems, the aforementioned approaches can not be generalized and they fail to generate reproducible results on unseen datasets. The application of pose estimation becomes a natural choice in our Visual Programming approach as it essentially pre-processes the video, focuses on specific human being (using the TRACK and CROP functionality) and then performs the pose classification. In our framework, we utilize the ‘Human model’ from MMPose [50] for keypoint detection in video frames. For example, as shown in Fig. 7 (Left) which depicts the person’s pose tagging in the video.
We then use these key points to classify the pose of a person (e.g. if a person is falling). This concept can be extended to hazard detection, street-crime detection, fight detection, yoga pose detection, and other applications.
For instance, consider a scenario in a video feed where a person is falling at the end Fig. 6. The user question, "What happens to the man at the end of the video?" will be decomposed into sub-tasks. It starts with triggering the TRIM step to filter out the relevant video frames only. This is followed by the POSE DETECT function applied to the preceding output frames from step 1. The output of POSE DETECT then acts as input to the succeeding CLASSIFY step, which detects that a man is falling.
4.3 Video Editing
Video editing is a crucial process in the post-production phase of film-making, television production, and other visual media industries [14]. It involves manipulating and rearranging video clips to create a coherent and engaging narrative or visual presentation. Video editing encompasses various tasks such as trimming, cutting, merging, and arranging video segments, as well as adding visual effects, transitions, and audio enhancements. In today’s digital age, video editing is typically performed using specialized software applications that offer a wide range of tools and features to facilitate the editing process.
Our approach directly applies to Video Editing by breaking any instruction down into a series of sequential steps. For this task we employ the use of the MERGE, CROP, TRIM, BGBLUR, and COLORPOP functions. Some concrete qualitative examples can be found in Fig. 1 and 7 (Middle).
Figure 7: Qualitative examples demonstrating three different use cases of the VURF framework including video editing, pose estimation and video question answering. The reasoning framework effectively understands the instructions and breaks down the problem into sub-tasks that are solved to reach the final outcome. Left: An example of Pose Detection that uses the visual program generated by VURF. Middle: A qualitative example of Video Editing which directly leverages VURF’s list of modules. Right: An example of VQA task decomposition by manageable sub-tasks by our framework. Best viewed in zoom.
5 Experiments and Results
Our experiments encompass both quantitative and qualitative evaluation of the proposed video reasoning framework. In the quantitative experiments, we assess the performance impact of the proposed video programming approach built on a pretrained model designed specifically for the Video Question Answering task. Additionally, we examine how the self-refinement approach affects the performance of the video programming approach, exploring the effects of altering the number of iterations on the model outputs. Finally, we present qualitative examples involving diverse tasks such as Pose Detection and Video Editing to highlight the efficacy of video programming in the context of video understanding tasks.
5.1 VQA Evaluation
In recent research endeavors, diverse approaches to video question answering (VQA) have been investigated, yielding promising outcomes through the utilization of large pre-trained vision-language models. Notably, the SeViLA framework [72], has emerged as a distinguished approach, effectively addressing important challenges in video processing. SeViLA integrates temporal keyframe localization and question answering by employing a unified image-language model (BLIP2)[35]. Demonstrating strong performance on VQA task, SeViLA achieves state-of-the-art accuracy across various VQA datasets, including STAR [63], NextQA [65], and Social IQ QA [74]. This paper incorporates SeViLA as the VQA module, showcasing improved performance in the zero-shot video question answering task with our VURF. The assessment is conducted on four benchmark datasets: STAR, NextQA, Social-IQ QA, and TrafficQA [66].
We conduct a comprehensive evaluation, commencing with the application of the base SeViLA zero-shot model on each dataset. Subsequently, we assess the effectiveness of the video programming approach on the same validation set. For the latter, a curated set of in-context examples is manually assembled for ICL, tailored to the specific characteristics of each dataset.
Several approaches, including InternVideo [62] and ViperGPT [55], have been proposed to address the challenge of Visual Question Answering alongside SeViLA. However, as illustrated in Table 1, VURF outperforms all three in zero-shot video question answering. Leveraging SeViLA as our VQA module provides a natural baseline for comparison. VURF offers a significant advantage over SeViLA and InternVideo by enabling reasoning through instructions rather than functioning as a black box, facilitating self-improvement. This capability proves advantageous over ViperGPT, which, despite employing the reasoning power of LLMs, is outperformed by VURF. The primary reason for this discrepancy is ViperGPT’s susceptibility to contextual hallucinations. We demonstrate that integrating self-refinement of in-context program-instruction pairs and incorporating a simple error correction module substantially enhances the performance of a visual programming approach.
While the concept of breaking down instructions or questions into simpler steps is generally beneficial, its effectiveness heavily relies on the functionality of the utilized functions. This limitation is evident in NextQA [65], particularly in categories involving temporal relations. Initially, we employed the GROUND functionality utilizing the TubeDETR [67] model. However, our observations revealed that when utilizing the GROUND/Localisation function, VURF does not surpass the baseline SeViLA model, primarily due to inaccuracies in the temporal localization functionality. To address this limitation and showcase the efficacy of the video programming approach, we manually annotated ground truths for the localization functionality in 30 randomly selected videos, resulting in improved accuracy (see Table 2). This underscores that with precise Temporal Localization functionality, VURF surpasses the performance of the base model. For the conducted testing, as depicted in Table 1, we deliberately refrain from utilizing the GROUND functionality to ensure fairness in the assessment. Although this omission slightly decreases the accuracy from 65.2% to 64.0%, it still exceeds that of SeViLA.
Finally, since the Social-IQ-2.0 dataset [74] contains the transcripts of the videos so we introduce a new function, ANALYSE, tailored for questions aimed at discerning the mood or tone of the video conversation. This function involves querying a Language Model (LLM) with the transcript, and, in some instances, the appended summary along with the posed question. Our analysis includes performance metrics for both scenarios i.e., employing transcripts and not using them (Table 3).
5.2 VQA Ablations
ViperGPT [55] diverges from VURF by not incorporating self-refinement mechanisms. This distinction is noteworthy as it highlights differing approaches to handling instruction comprehension within the context of VQA. Sole reliance on pre-trained language models without iterative refinement mechanisms makes the approach susceptible to contextual hallucinations. These hallucinations, stemming from the model’s inability to accurately contextualize instructions within the visual domain, can lead to erroneous outputs and decreased performance in tasks such as zero-shot video question answering. Our ablations (see Table 4) clearly demonstrate that integrating the self-refinement of in-context program-instruction pairs and incorporating a simple error correction module substantially enhances the performance of a visual programming approach.
Table 1: Performance of VURF on Video Question Answering (zero-shot) compared to other existing models. Each result is evaluated on the validation set of the corresponding dataset. The table shows the accuracies (in percentage) of the respective experiments.
Datasets NextQA STAR Social-IQ-2.0 TrafficQA
InternVideo [62] 50.2% 41.8% 30.1% 31.2%
ViperGPT [55] 60.0% 40.3% 37.8% 35.7%
SeViLA [72] 63.8% 44.3% 47.3% 39.1%
VURF 64.0% 47.2% 51.6% 43.5%
Table 2: Demostration of VURF without the influence of an imperfect module. We evaluate the performance of our Video Question Answering task on 30 randomly selected data points from the NextQA dataset by first using an imperfect temporal localization module (GROUND) and then replacing it by the actual ground truths of the localization queries. We find a noticeable increase in performance once a perfect Temporal Grounding model is introduced.
Dataset No Ground Truth Ground Truth
NextQA [65] 61.4% 65.2%
Table 3: Effect of using transcripts from the Social-IQ QA dataset. We evaluate the performance of our Video Question Answering task on Social-IQ QA dataset. The dataset contains transcripts so we conduct two experiments; with and without using transcripts. The results show an increase in performance (as expected) when the transcripts are utilized.
Dataset Baseline With transcripts
Social-IQ-2.0 [74] 48.1% 51.6%
5.3 Self-Refinement
  1. 1.
    In-Context Example Refinement: In examining the self-refinement process, we employ the NeXt-QA dataset by randomly sampling 50 videos for the test set. Additionally, we curate a set of 20 in-context examples for evaluation. The accuracy of the test set is initially calculated. Subsequently, these in-context examples undergo self-refinement. This approach initially presents instructions to an LLM for program generation without contextual influence. The generated program, along with the original program, is then provided to the LLM with a prompt to maintain the structure of the initial program while enhancing it using the non-contextual program. The refined in-context examples undergo multiple iterations of this process, and we report the accuracy on the test set with respect to the number of iterations of self-refinement applied to the in-context examples in Fig. 7(a).
  2. 2.
    Error Correction Evaluation: To assess the efficacy of automatic error correction, we randomly select 400 videos from the STAR dataset. The evaluation involves calculating the number of errors stemming from program invalidity. Given the ability to feed a program into the auto-correction module multiple times, we explore the impact of increasing these iterations on error occurrences within the 400 video-instruction pairs and report our evaluations in Fig. 7(b).
    (a) Accuracy plotted against a number of iterations of self-refinement We iteratively refine our in-context examples for a 50-video sample of NextQA and report our accuracy against the number of iterations of this self-refinement. The accuracy appears to quickly stabilize after 1 or 2 iterations.
    (b) Number of errors plotted against a number of iterations of self-refinement. The plot shows how the number of errors in the generated programs of a sample of 400 video-instruction pairs from the STAR dataset is affected by the number of iterations of the auto-error correction module.
    Figure 8: The impact of the self-refinement stage on quality of video programs.
    Table 4: Ablations demonstrating the effectiveness of the refinement pipeline. Results that do not use the error correction module, are such that if a syntactically incorrect program is generated it is considered as a wrong prediction.
    Datasets NextQA STAR Social-IQ-2.0 TrafficQA
    Accuracy w/o error correction & self refinement 47.8% 42.1% 45.5% 38.3%
    Accuracy w/o error correction 57.1% 44.9% 48.3% 41.1%
    Accuracy w/o self refinement 56.5% 43.1% 47.2% 40.2%
    Accuracy with self refinement & error correction 64.0% 47.2% 51.6% 43.5%
6 Conclusion
Our work expands the boundaries of Visual Programming by integrating it into the realm of video understanding, showcasing its efficacy in diverse applications such as Video Question Answering, Pose Estimation, and Multi-Video VQA. Additionally, we introduce a novel approach for enhancing the program generation process within the LLM through self-refinement, consequently elevating the efficacy of few-shot prompting to the LLM. Our approach not only broadens the scope of Visual Programming but also underscores the potential for continuous self-refinement to optimize the capabilities of LLMs for video reasoning tasks.
Appendix
7 Dataset Specific Prompts
We curated in-context examples for 4 different datasets (NeXTQA, STAR, Social-IQ, and TrafficQA). The prompts for two of the datasets are shown in Figure 9. We show 3 in-context examples for each prompt.
Figure 9: LEFT: LLM prompt for STAR dataset RIGHT: LLM prompt for TrafficQA dataset
8 Refinement Prompts
8.1 Error Correction
The error correction module queries the LLM two times, one to receive feedback on a program and the other to correct the program given the feedback. The prompts are shown in Figure 10. We display 2 in-context examples for each prompt.
Figure 10: LEFT: The prompt for the feedback generation of a given program. RIGHT: Given a feedback the correct program is generated by giving this prompt to the LLM.
8.2 Self-Refinement
The self-refinement module (which we pre-apply to refine our in-context examples) consists of 2 major queries to the LLM. One is for generating a context-free program to avoid hallucinations. Other is to convert the context-free program to a valid program. The prompts are shown in Figure 11
Figure 11: TOP: The prompt for the context-free generation. BOTTOM: The prompt for aligning a "perfect" program to a valid program
Figure 12: An example of the self refinement module. An initially generated prompt is refined by the LLM to produce a more complex and modular program.
Figure 13: A qualitative example of video-editing using VURF.
Figure 14: A step by step qualitative example of Video Question Answering using VURF.